OpenStudy (anonymous):

what is the domain of the following (use interval notation) : f(x)= 1/(3x+5)

OpenStudy (ash2326):

f(x)=1/(3x+5) 3x+5 can't be 0 i.e x cant be -5/3 so the domain is (-infinity, -5/3) U (-5/3, infinity) or Real-{-5/3}

OpenStudy (anonymous):

Anything that doesn't make (3x+5) = 0.

OpenStudy (anonymous):

also, how do I find the domain of \[f(x)=\sqrt{2x-7}\]

OpenStudy (anonymous):

I'm going to assume you're only dealing in real numbers, so you can't take the square-root of a negative. So (2x - 7) >= 0. Can you work it out from there?

OpenStudy (ash2326):

Swilly , inside the square root we should have a quantity greater than 0 2x-7>=0 2x>=7 x>=7

OpenStudy (anonymous):

7/2

OpenStudy (ash2326):

\[[3.5, \infty)\]

OpenStudy (ash2326):

sorry it should be 7/2 or 3.5

OpenStudy (anonymous):

ok thank you